The company had the following information for the month of January.

Cost of direct materials used in production $16,800
Direct labor $43,400
Factory overhead $28,000
Work in process inventory, January 1 $70,000
Work in process inventory, January 31 $74,200
Finished goods inventory, January 1 $29,400
Finished goods inventory, January 31 $33,600

For the month of January, determine (a) the costs of goods manufactured and (b) the cost of goods sold
Answer #1

Total manufacturing cost=Cost of direct materials used in production+Direct labor+Factory overhead

=16800+43400+28000

=$88200

Cost of goods manufactured=Beginning WIP+Total manufacturing cost-Ending WIP

=70,000+88200-74200

=$84000

Cost of goods sold=Beginning Finished goods inventory+Cost of goods manufactured-Ending Finished goods inventory

=29400+84000-33600

=$79800